OpenStudy (anonymous):

will award medal The slope of any line perpendicular segment AB to where A(0, -3) and B(4, 9) are two points on the line, is m = -3 -1/3 3

OpenStudy (gorv):

product of slope of perpendicular line=-1

OpenStudy (anonymous):

so -3?

OpenStudy (anonymous):

sorry, i dont understand

OpenStudy (gorv):

what is the slope of AB??

OpenStudy (gorv):

@emok9893

OpenStudy (anonymous):

i dont know what the formula is

OpenStudy (gorv):

u know how to find slope ??

OpenStudy (anonymous):

no

OpenStudy (gorv):

\[slope =\frac{ y2 -y1}{ x2-x1}\]

OpenStudy (gorv):

take one point as (x1,y1) and second as (x2,y2)

OpenStudy (anonymous):

how to do?

OpenStudy (anonymous):

im confused on how to set up the equation

OpenStudy (gorv):

\[\frac{ 9-(-3) }{ 4-0 }=??\]

OpenStudy (anonymous):

that equals 3

OpenStudy (gorv):

take one point as (x2,y2) let (4,9) is so x2=4 y2=9

OpenStudy (gorv):

yeah now slope of perpendicular = -1 /slope of the line

OpenStudy (anonymous):

-1/3

OpenStudy (anonymous):

thanks

OpenStudy (gorv):

welcome:)